OpenStudy (anonymous):

If y = (x + 3)2, and the value of x=5, what is the value?

OpenStudy (anonymous):

y=16

OpenStudy (anonymous):

y=(5+3)2 y=10+6 y=16 thats how i did it ... so i was right.. right?

OpenStudy (anonymous):

Yep, u r right!!

OpenStudy (anonymous):

thankkss

OpenStudy (anonymous):

no prob! :)